OpenStudy (anonymous):

Solve: 3x+2y+2z=-3 2x+3y+3z=-2 -3x-5y+z=-9

OpenStudy (anonymous):

Can you use matrices?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Well set it up like a matrix first like this

OpenStudy (anonymous):



OpenStudy (anonymous):

then reduce echelon form

OpenStudy (anonymous):

How would I do that on a TI-83 plus graphing calculator?

OpenStudy (anonymous):

You don't you do it by hand!!

OpenStudy (anonymous):

You won't be able to use your calculator in tests so you need to be able to do it by hand

OpenStudy (anonymous):

Ok but my son's math teacher is allowing calculator use but we don't know how to do it with this model of calculator. Do you know?

OpenStudy (anonymous):

Sorry I don't know how to do it with a calculator :/

OpenStudy (anonymous):

Thank you anyway

OpenStudy (whpalmer4):

Have you looked at the manual? I'm 100% certain that if the calculator has a function for doing this built in, it will describe how to do it.

OpenStudy (anonymous):

we found it online thanks! but we still need help on how to do it by hand :)

OpenStudy (whpalmer4):

Okay, now you're talking :-) \[3x+2y+2z=-3\]\[2x+3y+3z=-2\]\[-3x-5y+z=-9\] The idea is you're going to add and subtract various combinations of multiples of those equations to end up with some equations in fewer variables, eventually ending up with 1 equation in 1 variable, which after solving, you can back propagate the value through the equations until you have all of them.

OpenStudy (anonymous):

ok sounds good

OpenStudy (whpalmer4):

For example, adding the 1st and 3rd equations together would cause the x terms to disappear, giving you an equation in just y and z.

OpenStudy (anonymous):

thats is true...i did not think to do that!

OpenStudy (whpalmer4):

Multiplying the second equation by 3 and the third equation by 2 then adding them would similarly give you another equation in y and z.

OpenStudy (whpalmer4):

Now you have two equations in two unknowns, and I suspect that is not unfamiliar territory.

OpenStudy (anonymous):

no i can solve it from there...thank you so much for your help :)

OpenStudy (whpalmer4):

The matrix bit that RyanL mentioned is the way you want to do this for big systems, but in essence, what I described is what it does.

OpenStudy (anonymous):

thanks again!!

OpenStudy (whpalmer4):

Here's a brief write-up, if you're interested. Probably Khan Academy has a video on this as well. http://people.richland.edu/james/lecture/m116/matrices/matrices.html

OpenStudy (anonymous):

i appreciate it!